# Full quantum tomography of top quark decays

### v1

@@ Abstract
Quantum tomography in high-energy physics processes has usually been restricted to the spin degrees of freedom. We address the case of top quark decays $t\to Wb$, in which the angular momentum ($L$) and the spins of $W$ and $b$ are intertwined into a 54-dimensional $LWb$ density operator. The entanglement between $L$ and the $W$ or $b$ spin is large and could be determined for decays of single top quarks produced at the Large Hadron Collider, well above $5\sigma$ (statistical only) from the separability hypothesis. These would be the first entanglement measurements between orbital and spin angular momenta in high-energy physics. The method presented paves the way for similar measurements in other processes.
@@ 1 Introduction
Quantum tomography plays a crucial role in our understanding of particle physics processes. By reconstructing the density operator, quantum tomography provides a detailed snapshot of the quantum properties of a system, allowing us to gain insight into its behaviour. The full knowledge of the density operator enables the investigation of quantum entanglement patterns, and can also be used as a novel probe for new physics. For high-energy physics processes, previous literature has focused on the spin degrees of freedom of top quarks [1], $W$ [2] and $Z$ [3] bosons, or combinations of them [4, 5], which can be determined from multi-dimensional angular distributions of their decay products.
@@
Including the orbital angular momentum (hereafter referred to as $L$) is a non-trivial step further because momenta measurement prevents full direct access to its density operator $\rho_{L}$. High-energy physics experiments rely on the measurement of particle momenta, which projects over diagonal states. Specifically, when momenta $P$ are measured the resulting state is $\mathcal{P}\rho_{L}\mathcal{P}$, with $\mathcal{P}=|P\rangle\langle P|$, and different operators $\rho_{L}$ can lead to the same projection $\mathcal{P}\rho_{L}\mathcal{P}$. For the decay of top quarks and other two-body decays, the differential decay width is $d\Gamma\propto\langle\Omega|\rho_{L}|\Omega\rangle$, with $\Omega=(\theta,\phi)$ the flight direction of one of the decay products. $d\Gamma$ involves products of two spherical harmonics evaluated on the same arguments $(\theta,\phi)$, which are not orthogonal functions. Therefore, different operators $\rho_{L}$ may result in the same angular distribution. An extreme case is a scalar, e.g. the Higgs boson, where the decay distribution is isotropic despite $\rho_{L}$ contains non-zero entries $|l\,m\rangle\langle l^{\prime}\,m^{\prime}|$, up to $l,l^{\prime}=2$ in the decay to vector bosons and $l,l^{\prime}=1$ in the decay to fermions.
@@
In this Letter we consider a case of interest, the decay of top quarks $t\to Wb$, to address the determination of the 54-dimensional $LWb$ density operator that completely describes angular momentum in the top quark decay. We make a novel use of the complementarity between ‘canonical’ decay amplitudes, calculated using a fixed quantisation axis for all particles, and helicity amplitudes, in which the spin of the decay products is quantised in their direction of motion. While the former are the key tool to write down the $LWb$ density operator, the latter are very convenient to extract from multi-dimensional angular decay distributions the independent parameters that govern the top decay. The $LW$ and $Lb$ density operators are also of interest: they exhibit a large entanglement between the orbital and spin angular momentum parts, which can be experimentally determined, with a caveat that will be discussed. Previous experimental measurements of entanglement between orbital and spin angular momenta have been limited to photon pairs, see for example Refs. [6, 7].
2 Framework
@@
with $M$ the third spin component of $t$, $\lambda_{1}$ and $\lambda_{2}$ the helicities of $W$ and $b$, respectively, and $\lambda=\lambda_{1}-\lambda_{2}$. $D^{J}_{m^{\prime}m}$ are the Wigner functions and the ‘reduced amplitudes’ $a_{\lambda_{1}\lambda_{2}}$ are constants that do not depend on the angles, the only non-zero ones being $a_{1\frac{1}{2}}$, $a_{0\frac{1}{2}}$, $a_{0-\frac{1}{2}}$, $a_{-1-\frac{1}{2}}$. The combinations $\lambda_{1}=1$, $\lambda_{2}=-1/2$ and $\lambda_{1}=-1$, $\lambda_{2}=1/2$ are not allowed by angular momentum conservation, because the $L$ component in the direction of motion $\vec{p}$ vanishes. Performing a change of basis for the $b$ spinors and the $W$ polarisation vectors one can obtain the canonical amplitudes $A^{c}_{Ms_{1}s_{2}}(\theta,\phi)$, in which the $W$ and $b$ spins are quantised in the $\hat{z}$ axis, with respective eigenvalues $s_{1}$ and $s_{2}$. Expanding them in terms of spherical harmonics $Y_{l}^{m}$, and bearing in mind that $\langle\Omega|lm\rangle=Y_{l}^{m}(\Omega)$, the amplitudes $A_{Ms_{1}s_{2};lm}$ into $|lm\rangle$ eigenstates can be identified. The non-zero ones are
@@
The amplitudes in (3) are completely general, and the relation with $a_{\lambda_{1}\lambda_{2}}$ holds for any two-body decay of a spin-$1/2$ particle into massive spin-1 and spin-$1/2$ particles. For the top quark decay one can use the helicity amplitudes calculated in [9] for a generic interaction
@@
The masses of $t$, $W$ and $b$ are denoted as $m_{t}$, $M_{W}$ and $m_{b}$, as usual, and $E_{W}$, $E_{b}$ are the energies of $W$ and $b$, respectively, in the top quark rest frame. As a cross check, plugging (6) into (3) and using $\langle\Omega|lm\rangle=Y_{l}^{m}(\Omega)$, one recovers the canonical amplitudes $A^{c}_{Ms_{1}s_{2}}(\theta,\phi)$ obtained in [10] by direct calculation. Within the Standard Model (SM), and using $m_{t}=172.5$ GeV, $M_{W}=80.4$ GeV, $m_{b}=4.8$ GeV, one finds $(a_{1\frac{1}{2}},a_{0\frac{1}{2}},a_{0-\frac{1}{2}},a_{-1-\frac{1}{2}})=g(5.4,1.8,231,153)$ GeV, with $g$ the electroweak constant. The amplitudes with $\lambda_{2}=1/2$ are suppressed because the $tbW$ interaction is left-handed and the $b$ quark is much lighter than the top quark.
@@ 3 Density operators for top quark decay
Given a density operator
@@
with a sum over $M,M^{\prime}$. Taking the partial trace over the $b$ spin, $W$ spin or $L$ subspaces we obtain the density operators $\rho_{LW}$, $\rho_{Lb}$ and $\rho_{Wb}$, respectively. For these bipartite systems the entanglement can be characterised by the Peres-Horodecki [11, 12] criterion: for a density operator $\rho$ that describes a bipartite system $AB$, the positivity of the partial transpose over any subsystem, say $\rho^{T_{B}}$, is a necessary condition for separability. Conversely, a non-positive $\rho^{T_{B}}$ is a sufficient condition for entanglement. The amount of entanglement can be quantified by the negativity of the partial transpose on the $B$ subspace $\rho^{T_{B}}$ [13],
@@
where $\|X\|=\operatorname{tr}\sqrt{XX^{\dagger}}=\sum_{i}\sqrt{\lambda_{i}}$, where $\lambda_{i}$ are the (positive) eigenvalues of the matrix $XX^{\dagger}$. Because $\rho^{T_{B}}$ has unit trace, $N(\rho)$ equals the sum of the negative eigenvalues of $\rho^{T_{B}}$. (The result is the same when taking the partial transpose on subsystem $A$.) In the separable case $N(\rho)=0$. As an example, in the SM we have for the decay of top quarks in a $S_{3}=1/2$ eigenstate $N(\rho_{LW})=0.615$, $N(\rho_{Lb})=0.396$, $N(\rho_{Wb})=0.010$, i.e. all the pairs of subsystems are entangled when tracing on the remaining one. Note, however, that $\rho_{Wb}$ is separable in the $m_{b}=0$ limit, because in that case the $b$ quarks have left-handed helicity and there are only two non-zero $a_{\lambda_{1}\lambda_{2}}$, those with $\lambda_{2}=-1/2$. For unpolarised top quarks we have $N(\rho_{LW})=0.432$, $N(\rho_{Lb})=0.353$, $N(\rho_{Wb})=0$.
@@ 4 Parameter determination
Single top quarks are produced at the Large Hadron Collider (LHC) in the process $pp\to t\bar{b}j$ with a large polarisation in the direction of the spectator jet $j$. As shown in Ref. [14], this allows to determine $a_{\lambda_{1}\lambda_{2}}$, except for (i) a global normalisation that does not affect the density operators, and is set to unity; (ii) the relative phase $e^{i\alpha}$ between $\lambda_{2}=1/2$ and $\lambda_{2}=-1/2$ amplitudes, which has a marginal effect on the $LW$ and $Lb$ entanglement. In addition to the $(\theta,\phi)$ angles previously introduced for the top quark decay, to measure $W$ spin observables one has to consider the angles $(\theta^{*},\phi^{*})$ describing the orientation of the charged lepton momentum in the $W$ rest frame, with respect to a $(x^{\prime},y^{\prime},z^{\prime})$ system where the $\hat{z}^{\prime}$ axis is in the direction of $\vec{p}$ and the $\hat{y}^{\prime}$ angle lies in the $xy$ plane making an angle $\phi$ with the $\hat{y}$ axis.
The top quark polarisation, namely $P_{1-3}$, is determined by using the charged lepton as spin analyser [1]. We use forward-backward (FB) asymmetries on $\cos\theta_{\hat{n}}$, with $\theta_{\hat{n}}$ the angle between the charged lepton momentum in the top quark rest frame and an arbitrary axis $\hat{n}$. Their relation to the top quark polarisation $P_{\hat{n}}$ in that axis is
@@
$A_{\text{FB}}^{x^{\prime}z^{\prime}}=-\frac{1}{2\sqrt{2}}P_{3}\operatorname{Re}\left[a_{0\frac{1}{2}}a_{1\frac{1}{2}}^{*}-a_{-1-\frac{1}{2}}a_{0-\frac{1}{2}}^{*}\right]\,,$
$A_{\text{FB}}^{y^{\prime}z^{\prime}}=-\frac{1}{2\sqrt{2}}P_{3}\operatorname{Re}\left[a_{0\frac{1}{2}}a_{1\frac{1}{2}}^{*}-a_{-1-\frac{1}{2}}a_{0-\frac{1}{2}}^{*}\right]\,.$ (15)
Finally, the relative phase between $\lambda_{2}=1/2$ and $\lambda_{2}=-1/2$ amplitudes cannot be measured as long as the $b$ quark polarisation is not currently accessible in the ATLAS and CMS experiments. However, its impact is marginal because $\lambda_{2}=1/2$ amplitudes are quite suppressed. For top quarks in a $S_{3}=1/2$ eigenstate, and using the SM values for $a_{\lambda_{1}\lambda_{2}}$ but inserting an arbitrary relative phase $e^{i\alpha}$ in $a_{1\frac{1}{2}}$ and $a_{0\frac{1}{2}}$, $N(\rho_{LW})$ lies in the range $[0.614,0.628]$ and $N(\rho_{Lb})$ in the range $[0.396,0.411]$. The minimum entanglement is found for the SM value of the phase, $\alpha=0$. This variation, of the order of the expected statistical uncertainty, is not relevant for the determination of the $LW$ and $Lb$ entanglement.

### v2

@@ Abstract
Quantum tomography in high-energy physics processes has usually been restricted to the spin degrees of freedom. We address the case of top quark decays $t\to Wb$, in which the orbital angular momentum ($L$) and the spins of $W$ and $b$ are intertwined into a 54-dimensional $LWb$ density operator. The entanglement between $L$ and the $W$ or $b$ spin is large and could be determined for decays of single top quarks produced at the Large Hadron Collider with Run 2 data. With the foreseen statistical and systematic uncertainties, the significance is well above $5\sigma$ from the separability hypothesis for $L$-$W$ entanglement, and $3.2\sigma$ for $L$-$b$. These would be the first entanglement measurements between orbital and spin angular momenta in high-energy physics. Likewise, the genuine tripartite entanglement between $L$ and the two spins could be established with more than $5\sigma$. The method presented paves the way for similar measurements in other processes.
@@ 1 Introduction
Quantum tomography plays a crucial role in our understanding of particle physics processes. By reconstructing the density operator, quantum tomography provides a detailed snapshot of the quantum properties of a system, allowing us to gain insight into its behaviour. The full knowledge of the density operator enables the investigation of quantum entanglement patterns, and can also be used as a novel probe for new physics. For high-energy physics processes, previous literature has focused on the spin degrees of freedom of top quarks [1], $W$ [2] and $Z$ [3] bosons, or combinations of them [4, 5, 6], which can be determined from multi-dimensional angular distributions of their decay products. Likewise, spin entanglement tests have been proposed for top quarks [7, 8, 9, 10, 11, 12, 13, 14], muons [15], $\tau$ leptons [16], weak bosons [17, 18, 19, 20, 5, 21, 22], and particles of different spin [23, 24, 25]. Spin-spin entanglement has been experimentally probed with top quark pairs by the ATLAS [26] and CMS [27] Collaborations.
@@
Including the orbital angular momentum (hereafter referred to as $L$) is a non-trivial step further because momenta measurement prevents full direct access to its density operator $\rho_{L}$. High-energy physics experiments rely on the measurement of particle momenta, which projects over diagonal states. Specifically, when momenta $P$ are measured the resulting state is $\mathcal{P}\rho_{L}\mathcal{P}$, with $\mathcal{P}=|P\rangle\langle P|$, and different operators $\rho_{L}$ can lead to the same projection $\mathcal{P}\rho_{L}\mathcal{P}$. For the decay of top quarks and other two-body decays, the differential decay width is $d\Gamma\propto\langle\Omega|\rho_{L}|\Omega\rangle$, with $\Omega=(\theta,\phi)$ the flight direction of one of the decay products. $d\Gamma$ involves products of two spherical harmonics $\Pi_{ll^{\prime}}^{mm^{\prime}}(\theta,\phi)\equiv Y_{l}^{m}(\theta,\phi)Y_{l^{\prime}}^{m^{\prime}}(\theta,\phi)^{*}$ evaluated on the same arguments. These products $\Pi_{ll^{\prime}}^{mm^{\prime}}(\theta,\phi)$ are not orthogonal functions, therefore, different operators $\rho_{L}$ may result in the same angular distribution. An extreme case is a scalar, e.g. the Higgs boson, where the decay distribution is isotropic despite $\rho_{L}$ contains non-zero entries $|l\,m\rangle\langle l^{\prime}\,m^{\prime}|$, up to $l,l^{\prime}=2$ in the decay to vector bosons and $l,l^{\prime}=1$ in the decay to fermions.
@@
In this Letter we consider a case of interest, the decay of top quarks $t\to Wb$, to address the determination of the 54-dimensional $LWb$ density operator that completely describes angular momentum in the top quark decay. We make a novel use of the complementarity between ‘canonical’ decay amplitudes, calculated using a fixed quantisation axis for all particles, and helicity amplitudes, in which the spin of the decay products is quantised in their direction of motion. While the former are the key tool to write down the $LWb$ density operator, the latter are very convenient to extract from multi-dimensional angular decay distributions the independent parameters that govern the top decay. The $LW$ and $Lb$ density operators are also of interest: they exhibit a large entanglement between the orbital and spin angular momentum parts, which can be experimentally determined, with a caveat that will be discussed. Previous experimental measurements of entanglement between orbital and spin angular momenta have been limited to photon pairs, see for example Refs. [28, 29]. The genuine tripartite entanglement between $L$ and the two spins is also large, and could be experimentally established.
2 Top quark decay amplitudes
@@
with $M$ the third spin component of $t$, $\lambda_{1}$ and $\lambda_{2}$ the helicities of $W$ and $b$, respectively, and $\lambda=\lambda_{1}-\lambda_{2}$. $D^{J}_{m^{\prime}m}$ are the Wigner functions and the ‘reduced amplitudes’ $a_{\lambda_{1}\lambda_{2}}$ are constants that do not depend on the angles, the only non-zero ones being $a_{1\frac{1}{2}}$, $a_{0\frac{1}{2}}$, $a_{0-\frac{1}{2}}$, $a_{-1-\frac{1}{2}}$. The combinations $\lambda_{1}=1$, $\lambda_{2}=-1/2$ and $\lambda_{1}=-1$, $\lambda_{2}=1/2$ are not allowed by angular momentum conservation, because the $L$ component in the direction of motion $\vec{p}$ vanishes. Performing a change of basis for the $b$ spinors and the $W$ polarisation vectors one can obtain the canonical amplitudes $A^{c}_{Ms_{1}s_{2}}(\theta,\phi)$, in which the $W$ and $b$ spins are quantised in the $\hat{z}$ axis, with respective eigenvalues $s_{1}$ and $s_{2}$. Expanding them in terms of spherical harmonics $Y_{l}^{m}$, and bearing in mind that $\langle\Omega|lm\rangle=Y_{l}^{m}(\Omega)$, the amplitudes $A_{Ms_{1}s_{2};lm}$ for decay to $|s_{1}s_{2}lm\rangle$ eigenstates can be identified. The non-zero ones are
@@
We note that the dependence on angles $(\theta,\phi)$ is no longer present in the amplitudes (3) for decay to $|s_{1}s_{2}lm\rangle$ eigenstates. These amplitudes are completely general, and the relation with $a_{\lambda_{1}\lambda_{2}}$ holds for any two-body decay of a spin-$1/2$ particle into massive spin-1 and spin-$1/2$ particles. For the top quark decay one can use the helicity amplitudes calculated in [24] for a generic interaction
@@
The masses of $t$, $W$ and $b$ are denoted as $m_{t}$, $M_{W}$ and $m_{b}$, as usual, and $E_{W}$, $E_{b}$ are the energies of $W$ and $b$, respectively, in the top quark rest frame. As a cross check, plugging (6) into (3) and using $\langle\Omega|lm\rangle=Y_{l}^{m}(\Omega)$, one recovers the canonical amplitudes $A_{Ms_{1}s_{2}}(\theta,\phi)$ obtained in [25] by direct calculation. Within the Standard Model (SM), and using $m_{t}=172.5$ GeV, $M_{W}=80.4$ GeV, $m_{b}=4.8$ GeV, one finds $(a_{1\frac{1}{2}},a_{0\frac{1}{2}},a_{0-\frac{1}{2}},a_{-1-\frac{1}{2}})=g(5.4,1.8,231,153)$ GeV, with $g$ the electroweak constant. The amplitudes with $\lambda_{2}=1/2$ are suppressed because the $tbW$ interaction is left-handed and the $b$ quark is much lighter than the top quark.
@@ 3 Density operators for top quark decay
Given a density operator parameterised as
@@
with a sum over $M,M^{\prime}$. Taking the partial trace over the $b$ spin, $W$ spin or $L$ subspaces we obtain the density operators $\rho_{LW}$, $\rho_{Lb}$ and $\rho_{Wb}$, respectively, with matrix elements
$(\rho_{LW})_{s_{1};lm}^{s_{1}^{\prime};l^{\prime}m^{\prime}}=(\rho_{t})_{MM^{\prime}}A_{Ms_{1}s_{2};lm}A_{M^{\prime}s_{1}^{\prime}s_{2};l^{\prime}m^{\prime}}\,,$
$(\rho_{Lb})_{s_{2};lm}^{s_{2}^{\prime};l^{\prime}m^{\prime}}=(\rho_{t})_{MM^{\prime}}A_{Ms_{1}s_{2};lm}A_{M^{\prime}s_{1}s_{2}^{\prime};l^{\prime}m^{\prime}}^{*}\,,$
$(\rho_{Wb})_{s_{1}s_{2}}^{s_{1}^{\prime}s_{2}^{\prime}}=(\rho_{t})_{MM^{\prime}}A_{Ms_{1}s_{2};lm}A_{M^{\prime}s_{1}^{\prime}s_{2}^{\prime};lm}^{*}\,,$ (10)
with implicit sums over repeated indices. The operators $\rho_{LW}$ and $\rho_{Lb}$ have an obvious interpretation: they correspond to marginalising the $b$ quark or $W$ boson spin degrees of freedom. The interpretation of $\rho_{Wb}$ is less straightforward. Since in $\rho_{LWb}$ there is no angular dependence and the $L$ degrees of freedom are marginalised, this operator describes the $Wb$ spin degrees of freedom integrated over all top quark decay phase space. An independent computation of $\rho_{Wb}$ using the canonical decay amplitudes $A^{c}_{Ms_{1}s_{2}}$, followed by integration over $(\theta,\phi)$, renders exactly the same result for $\rho_{Wb}$.
For bipartite systems the entanglement can be characterised by the Peres-Horodecki [31, 32] criterion: for a density operator $\rho$ that describes a bipartite system $AB$, the positivity of the partial transpose over any subsystem, say $\rho^{T_{B}}$, is a necessary condition for separability. Conversely, a non-positive $\rho^{T_{B}}$ is a sufficient condition for entanglement. The amount of entanglement can be quantified by the negativity of the partial transpose on the $B$ subspace $\rho^{T_{B}}$ [33],
@@
where $\|X\|=\operatorname{tr}\sqrt{XX^{\dagger}}=\sum_{i}\sqrt{\lambda_{i}}$, where $\lambda_{i}$ are the (positive) eigenvalues of the matrix $XX^{\dagger}$. Because $\rho^{T_{B}}$ has unit trace, $N(\rho)$ equals the sum of the negative eigenvalues of $\rho^{T_{B}}$. (The result is the same when taking the partial transpose on subsystem $A$.) In the separable case $N(\rho)=0$. As an example, in the SM we have for the decay of top quarks in a $S_{3}=1/2$ eigenstate $N(\rho_{LW})=0.615$, $N(\rho_{Lb})=0.396$, $N(\rho_{Wb})=0.010$, i.e. all the pairs of subsystems are entangled when tracing on the remaining one. Note, however, that $\rho_{Wb}$ is separable in the $m_{b}=0$ limit, because in that case the $b$ quarks have left-handed helicity and there are only two non-zero $a_{\lambda_{1}\lambda_{2}}$, those with $\lambda_{2}=-1/2$. For unpolarised top quarks we have $N(\rho_{LW})=0.432$, $N(\rho_{Lb})=0.353$, $N(\rho_{Wb})=0$. The $W$-$b$ entanglement, which is zero or nearly zero in all situations, is ignored from now on.
The genuine tripartite entanglement can also be established. Namely, the fact that the state is not separable under any bipartition of the Hilbert space $\mathcal{H}_{L}\otimes\mathcal{H}_{W}\otimes\mathcal{H}_{b}$ in which $\rho_{LWb}$ acts. (For easier notation we label the Hilbert spaces for the $W$ and $b$ spins as $\mathcal{H}_{W}$ and $\mathcal{H}_{b}$, respectively.) For example, the $L$-$(Wb)$ entanglement can be tested by using the Peres-Horodecki sufficient condition with $A=\mathcal{H}_{L}$, $B=\mathcal{H}_{W}\otimes\mathcal{H}_{b}$, and the amount of entanglement can be quantified by the negativity $N$ in (11). Likewise, the $W$-$(Lb)$ and $b$-$(LW)$ entanglement can be established. For the decay of top quarks in a $S_{3}=1/2$ eigenstate we have $N(\rho_{L[Wb]})=1.65$, $N(\rho_{W[Lb]})=0.80$, $N(\rho_{b[LW]})=0.50$ for $L$-$(Wb)$, $W$-$(Lb)$ and $b$-$(LW)$ entanglement, respectively. For unpolarised top quarks we have $N=1.24,0,59,0.49$.
@@ 4 Parameter determination
Single top quarks are produced at the Large Hadron Collider (LHC) in the process $pp\to t\bar{b}j$ with a large polarisation in the direction of the spectator jet $j$. As shown in Ref. [34], this allows to determine $a_{\lambda_{1}\lambda_{2}}$, except for (i) a global normalisation that does not affect the density operators, and is set to unity; (ii) the relative phase $e^{i\alpha}$ between $\lambda_{2}=1/2$ and $\lambda_{2}=-1/2$ amplitudes, which has a marginal effect on the entanglement because the former are quite suppressed. In addition to the $(\theta,\phi)$ angles previously introduced for the top quark decay, to measure $W$ spin observables one has to consider the angles $(\theta^{*},\phi^{*})$ describing the orientation of the charged lepton momentum in the $W$ rest frame, with respect to a $(x^{\prime},y^{\prime},z^{\prime})$ system where the $\hat{z}^{\prime}$ axis is in the direction of $\vec{p}$ and the $\hat{y}^{\prime}$ angle lies in the $xy$ plane making an angle $\phi$ with the $\hat{y}$ axis.
The top quark polarisation, namely $P_{1-3}$ in Eq. (8), is determined by using the charged lepton as spin analyser [1]. We use forward-backward (FB) asymmetries on $\cos\theta_{\hat{n}}$, with $\theta_{\hat{n}}$ the angle between the charged lepton momentum in the top quark rest frame and an arbitrary axis $\hat{n}$. Their relation to the top quark polarisation $P_{\hat{n}}$ in that axis is
@@
$A_{\text{FB}}^{x^{\prime}z^{\prime}}=-\frac{3\pi}{8\sqrt{2}}P_{3}\operatorname{Re}\left[a_{0\frac{1}{2}}a_{1\frac{1}{2}}^{*}-a_{-1-\frac{1}{2}}a_{0-\frac{1}{2}}^{*}\right]\,,$
$A_{\text{FB}}^{y^{\prime}z^{\prime}}=-\frac{3\pi}{8\sqrt{2}}P_{3}\operatorname{Re}\left[a_{0\frac{1}{2}}a_{1\frac{1}{2}}^{*}-a_{-1-\frac{1}{2}}a_{0-\frac{1}{2}}^{*}\right]\,.$ (16)
Finally, the relative phase between $\lambda_{2}=1/2$ and $\lambda_{2}=-1/2$ amplitudes cannot be measured as long as the $b$ quark polarisation is not currently measured for top quark decays in the ATLAS and CMS experiments. However, its impact on entanglement is marginal because $\lambda_{2}=1/2$ amplitudes are quite suppressed. [^1] In scenarios beyond the SM where $\lambda_{2}=1/2$ amplitudes are not supressed (e.g. in the presence of $\bar{b}_{R}\sigma^{\mu\nu}t_{L}$ or $\bar{b}_{R}\gamma^{\mu}t_{R}$ interactions) this would not be the case. The presence of such interactions could easily be detected by a non-vanishing value of $F_{+}$, and would also induce deviations in the asymmetries in (15) and (16) and polarisations $P_{1}$, $P_{2}$ from their SM prediction [34]. For top quarks in a $S_{3}=1/2$ eigenstate, and using the SM values for $a_{\lambda_{1}\lambda_{2}}$ but inserting an arbitrary relative phase $e^{i\alpha}$ in $a_{1\frac{1}{2}}$ and $a_{0\frac{1}{2}}$, $N(\rho_{LW})$ lies in the range $[0.614,0.628]$ and $N(\rho_{Lb})$ in the range $[0.396,0.411]$. The minimum entanglement is found for the SM value of the phase, $\alpha=0$. This variation, of the order of the expected statistical uncertainty, is not relevant for the determination of the $L$-$W$ and $L$-$b$ entanglement. Similarly, the variation in the entanglement between different bipartitions of $\mathcal{H}_{L}\otimes\mathcal{H}_{W}\otimes\mathcal{H}_{b}$, for different values of $\alpha$, is unimportant. The maximum difference with respect to the SM is obtained for $\alpha=\pi$: $+0.025$ for $N(\rho_{L[Wb]})$ and $-10^{-3}$ for $N(\rho_{b[LW]})$, with $O(10^{-6})$ differences for $N(\rho_{W[Lb]})$.